### v1

@@ 2.3 Arrival Time Model: Hyperbolic Wavefront
Figure 2: Dependence of the optimal timing parameter $C$ on the slant distance to the shower maximum, $D_{X_{\mathrm{max}}}$. Each point represents an individual CoREAS simulation, where the signal times were determined via the Hilbert envelope maximum of the total electric field trace, colored by the root-mean-square (RMS) of the timing residuals for the best fit. The solid black line shows the fit using a 3rd-degree polynomial (Tab. 1), with the shaded region indicating the root-mean-square (RMS) deviations from the fit. The dashed gray line indicates the fixed value used by LOPES, $C=25200~\mathrm{g/cm}^{2}$).
@@ 2.4 Particle Model: NKG Function
The full particle model is described in Appendix A.
@@ 4.3.1 $E_{\text{rad}}$ reconstruction performance
Figure 4: Performance of the radiation energy $E_{\text{rad}}$ reconstruction on CoREAS simulations. The left side displays both the reconstructed values compared to the CoREAS truth for all 390 successful reconstructions, as well as the pull from the reconstruction. The right side of the plot shows a histogram of the reconstruction errors at the top and how the reconstruction uncertainties correlate with reconstruction error.
@@ 4.3.2 $X_{\text{max}}$ reconstruction performance
Figure 5: Performance of the $X_{\text{max}}$ reconstruction on CoREAS simulations. (Left) Reconstructed vs. true $X_{\text{max}}$ and pull distribution. (Right) Error distribution and uncertainty correlation.
Analogously, the performance on the $X_{\text{max}}$ is shown in Fig. 5. We achieve an overall precision on the $X_{\text{max}}$ of $25.4\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$ with a bias of $0.3\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$. The Pull $\sigma$ of 1.52 indicates that the uncertainties are slightly underestimated. We observe that especially for larger shower maxima, where the footprints tend to extend only over $100\text{\,}\mathrm{m}$\$, the reconstruction predicts lower shower maxima. Since both the fluence and the timing parameterization are less accurate for large $X_{\text{max}}$, this behaviour is somewhat expected.
Additionally to the combined timing-fluence model, we also tested an only fluence-based reconstruction. This was to test how much the timing information contributes to the fit. For this, we prepared the data exactly as for the timing-fluence case, but discarded timing data and the time component of the forward model. With that approach, we achieved a resolution in $X_{\text{max}}$ of $36\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$, meaning the timing information improved the resolution substantially, as further discussed in Appendix 11.
@@ 4.3.3 Shower geometry performance
The fluence-based fit, in which timing data was discarded, achieved a similar resolution in azimuth as the combined fit, but only a $1.5\text{\,}\mathrm{\SIUnitSymbolDegree}$ resolution on the zenith angle. This further shows how the timing data is beneficial to the overall reconstruction (see Appendix 11).
@@ 5 Application to LOFAR data
Post-reconstruction we apply the calculated pull factors from the simulation results (Fig. 4, 5) to the reconstruction uncertainties (i.e. we scale the uncertainties by the pull factor) to avoid underestimating them.
@@ 5.1 Reconstruction quality
Furthermore, we show the event-wise comparison in reconstructed $X_{\text{max}}$ and $E_{\text{CR}}$ in Fig. 9. The $X_{\text{max}}$ distribution on the right shows a weighted bias of $0.55\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$ and a standard deviation of $52.4\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$. Assuming the statistical uncertainties of the two methods are uncorrelated, the expected standard deviation of the difference distribution based on the respective resolutions would be $\sigma_{\Delta}=\sqrt{\sigma_{\text{IFT}}^{2}+\sigma_{\text{2021}}^{2}}\approx$33\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$\$. Given that the standard deviation in Fig. 9 is larger than that, the precision of either method might be overestimated. Without independent measurements it is impossible to determine, which. This is also reflected in the Pull $\sigma$ of 2.78 of this distribution. The small bias between the two however does not indicate a fundamental systematic difference between them.
@@ Acknowledgments
Appendix A Particle Model

### v2

@@ 2.3 Arrival Time Model: Hyperbolic Wavefront
Figure 2: Dependence of the optimal timing parameter $C$ on the slant distance to the shower maximum, $D_{X_{\mathrm{max}}}$. Each point represents an individual CoREAS simulation, where the signal times were determined via the Hilbert envelope maximum of the total electric field trace, colored by the root-mean-square (RMS) of the timing residuals for the best fit. The solid black line shows the fit using a 3rd-degree polynomial (Tab. 1), with the shaded region indicating the root-mean-square (RMS) deviations from the fit. The dashed gray line indicates the fixed value used by LOPES, $C=25200~\mathrm{g/cm}^{2}$.
@@ 2.4 Particle Model: NKG Function
The full particle model is described in Appendix C.
@@ 4.3 Reconstruction performance
We find that especially for higher energy events, the posteriors resulting from the reconstruction are too narrow. For lower energies, noise dominates the data, which is reflected in broader posterior distributions. For higher energy events, systematic data-model differences dominate, resulting lower uncertainties. Activating the fluence correlated field could be a solution to that in the future. For now, we perform a calibration of uncertainties based on the reconstructed radiation energy. The full calibration procedure is detailed in Appendix B. This calibration has been applied to all of the results that are discussed in the following sections.
We investigated, whether this underestimation of uncertainties is inherently due to the usage of VI, which in some cases can lead to too narrow posteriors. Running the reconstruction again using a No U-Turn Sampler (NUTS) from the NIFTY framework revealed no significant change in posterior width, indicating that the under-coverage stems from the model and not the inference engine.
@@ 4.3.1 $E_{\text{rad}}$ reconstruction performance
Figure 4: Performance of the radiation energy $E_{\text{rad}}$ reconstruction on CoREAS simulations. The left side displays both the reconstructed values compared to the CoREAS truth for all 390 successful reconstructions, as well as the pull from the reconstruction. The color-coding of the scatter points is based on how many stations were considered "triggered" for this specific event. The right side of the plot shows a histogram of the reconstruction errors at the top and how the reconstruction uncertainties correlate with reconstruction error.
@@ 4.3.2 $X_{\text{max}}$ reconstruction performance
Figure 5: Performance of the $X_{\text{max}}$ reconstruction on CoREAS simulations. (Left) Reconstructed vs. true $X_{\text{max}}$ and pull distribution. The color-coding of the scatter points is based on how many stations were considered "triggered" for this specific event. (Right) Error distribution and uncertainty correlation.
Analogously, the performance on the $X_{\text{max}}$ is shown in Fig. 5. We achieve an overall precision on the $X_{\text{max}}$ of $25.4\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$ with a bias of $0.3\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$. The Pull $\sigma$ post-calibration of uncertainties is 0.94. Before calibration, the Pull $\sigma$ was 1.52.
We observe that especially for larger shower maxima, where the footprints tend to extend only over $100\text{\,}\mathrm{m}$\$, the reconstruction predicts lower shower maxima. Since both the fluence and the timing parameterization are less accurate for large $X_{\text{max}}$, this behaviour is somewhat expected.
Additionally to the combined timing-fluence model, we also tested an only fluence-based reconstruction. This was to test how much the timing information contributes to the fit. For this, we prepared the data exactly as for the timing-fluence case, but discarded timing data and the time component of the forward model. With that approach, we achieved a resolution in $X_{\text{max}}$ of $36\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$, meaning the timing information improved the resolution substantially, as further discussed in Appendix A.
@@ 4.3.3 Shower geometry performance
The fluence-based fit, in which timing data was discarded, achieved a similar resolution in azimuth as the combined fit, but only a $1.5\text{\,}\mathrm{\SIUnitSymbolDegree}$ resolution on the zenith angle. This further shows how the timing data is beneficial to the overall reconstruction (see Appendix A).
@@ 5 Application to LOFAR data
Post-reconstruction we apply the simulation-based calibration described in Appendix B to the reconstruction uncertainties to avoid underestimating them.
@@ 5.1 Reconstruction quality
Furthermore, we show the event-wise comparison in reconstructed $X_{\text{max}}$ and $E_{\text{CR}}$ in Fig. 9. The $X_{\text{max}}$ distribution on the right shows no bias and a standard deviation of $52.4\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$. Assuming the statistical uncertainties of the two methods are uncorrelated, the expected standard deviation of the difference distribution based on the respective resolutions would be $\sigma_{\Delta}=\sqrt{\sigma_{\text{IFT}}^{2}+\sigma_{\text{2021}}^{2}}\approx$33\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$\$. Upon further investigation of the differences in $X_{\text{max}}$ between the two methods, we notice that this large scatter is primarily outlier-driven, as the difference distribution is slightly asymmetric with a non-Gaussian tail, indicating a few strong outliers where the IFT method predicts much larger $X_{\text{max}}$ values than the legacy method. To account for this, we recalculated the scatter using a robust standard deviation estimator: the $68\%$ interquantile range, defined as $\sigma_{\mathrm{rob}}=(P_{84}-P_{16})/2$. This metric isolates the true $1\sigma$ width of the dense core of the data while ignoring the extreme outlier tails. Using this metric, we retrieve a scatter of only $35\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$, however we also find that the median difference between the two methods is $9.7\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-2}$ using this metric. Nevertheless, for the absolute majority of reconstructions the methods agree within their nominal precisions for $X_{\text{max}}$. The strong outliers warrant further investigation for a future composition study.
@@ Acknowledgments
Figure 12: Pull standard deviation of the $E_{\text{rad}}$ for different shower parameters: SNR (maximum fluence SNR across the event), number of triggered stations, reconstructed radiation energy and zenith angle. The original Pull $\sigma$ are compared to a simple global correction and a correction based on a cubic fit to the Pull $\sigma$ as function of reconstructed radiation energy.
Figure 13: Pull standard deviation of the $X_{\text{max}}$ for different shower parameters: SNR (maximum fluence SNR across the event), number of triggered stations, reconstructed radiation energy and zenith angle. The original Pull $\sigma$ are compared to a simple global correction and a correction based on a cubic fit to the Pull $\sigma$ as function of reconstructed radiation energy.
Appendix A Impact of added timing information
Appendix B Calibration of uncertainties
To account for the underestimation of uncertainties, we calibrate based on the Pull $\sigma$ of our simulation set. Fig. 12 shows the Pull $\sigma$ as a function of shower parameters for the radiation energy.
The reconstruction appears overly confident for higher energy events, likely because for lower energies the data are more dominated by noise than by systematic differences from the forward model. We model the Pull $\sigma$ as a function of reconstructed radiation energy via a third degree polynomial fit. The posterior distribution of the radiation energy is then scaled by this function. Fig. 12 also shows how a global correction by the overall Pull $\sigma$ would perform. Overall we find the $E_{\text{rad}}^{\text{reco}}$-based correction yields a better coverage.
The $X_{\text{max}}$ reconstruction in Fig. 5 also indicates undercoverage. As for the radiation energy, we find that the Pull increases as a function of energy, so we calibrate based on reconstructed radiation energy. Fig. 13 shows the Pull dependence for the $X_{\text{max}}$ as function of shower parameters, the cubic fit used for the calibration and the comparison with a simple global correction.
We note that there seems to be an additional dependence on zenith angle, as the bottom right subfigure of Fig. 13 indicates. We attempted also a two-dimensional calibration based on both radiation energy and zenith angle, however we found that the resulting overall coverage in $X_{\text{max}}$ was worse than for a one-dimensional calibration.
We suspect that, especially for the radiation energy reconstruction, the usage of correlated fields in the fluence forward model would provide a better coverage due to the added degrees of freedom in the model. In the future, the effect of the fluence correlated field on $X_{\text{max}}$ could be further investigated, possibly resulting in a similar reconstruction precision and better coverage without the need for a simulation-based calibration.
Appendix C Particle Model